A family is finishing a room in their basement. The room is 20 feet long and 14 feet wide. There is a 2'8" - wide door opening in one wall. How many linear feet of wall base are needed? Note that no wall base is needed along the open door opening.
Answer by math_tutor2020(3816)


For now, let's say there isn't a door. If so, then the amount of wall base needed is exactly the perimeter of the rectangle with L = 20 as the length and W = 14 as the width.

Apply the perimeter of a rectangle formula with those values
P = 2*(L+W)
P = 2*(20+14)
P = 2*(34)
P = 68

The perimeter of the rectangle is 68 feet. If we didn't have a door, then this would be the answer. This is assuming you don't have any waste/errors.

However, we do have to account for the door.

So we'll subtract off the 2'8" like so
68' - 2'8"
68 ft - (2 ft, 8 in)
68 ft - (2 ft + 8 in)
68 ft - 2 ft - 8 in
66 ft - 8 in
65 ft + 1 ft - 8 in
65 ft + 12 in - 8 in
65 ft + 4 in
65 ft, 4 in
65' 4"

Note how I broke up the 66 ft into 65 ft + 1 ft. Afterward, that 1 ft is turned into 12 inches. This allows us to apply the subtraction later on (this is an example of "borrowing" in subtraction).

After accounting for the door, we'll need exactly 65'4" of wall base assuming that we don't make any mistakes. Though for realistic purposes, it's better to overshoot your goal to allow better flexibility.

Answer: 65' 4"